# FreeBird.jl: An Extensible Toolbox for Simulating Interfacial Phase Equilibria

### v1

@@ Abstract
[…]
@@ 1 Introduction
Because high-resolution in situ probes remain scarce, 10, 11 theoretical and computational approaches have become indispensable for mapping the configurational landscape of interfaces and guiding materials design. 12, 13 Modeling an interface under working conditions is essentially a non-equilibrium surface problem whose dynamics depend on the poorly constrained, synthesis-specific starting structure. 14, 15 To address this, researchers typically begin by establishing an equilibrium thermodynamic baseline, 16, 17 which entails a series of decisions: the level of energy evaluation (first-principles versus classical force fields) 18, 19 and the depth of configurational sampling required to obtain entropies and free energies (harmonic, quasi-harmonic, or fully anharmonic treatments; atomistic versus lattice representations). 20, 21, 22
@@
Most realistic surface problems are too complex for exhaustive first-principles treatment, forcing researchers to navigate an accuracy–sampling trade-off. 16, 23 Ab initio thermodynamics, the workhorse for predicting surface phase diagrams, typically evaluates only tens to thousands of candidate structures within a practical computing budget, with selections often guided by chemical intuition. 16, 17 Structure-search algorithms such as particle-swarm, 24 Bayesian, 25, 26 and genetic 27, 28 optimizers expand the candidate pool but still assume a sharply truncated configuration space in which atoms vibrate harmonically about fixed lattice sites. 20, 22 This harmonic perspective understates the structural and compositional agility of catalytic surfaces; in many cases, comprehensive sampling with well-parameterized classical potentials can reveal low-energy motifs that sparse, high-accuracy DFT surveys may miss. 18, 19, 29, 30
@@
Although progress has been made toward systematically improvable thermodynamic sampling of surface phases, 31, 32, 29 the field still lacks a fast, reproducible benchmarking platform dedicated to interfacial structure prediction, applicable to solid–vacuum/gas systems as well as defects at solid–solid and solid–liquid boundaries. We therefore introduce FreeBird.jl, an open-source Julia package featuring a highly parallel, modular architecture that allows users to interchange system resolutions, energy models, and Monte Carlo–style algorithms without rewriting code. Distinct from existing packages, 33, 34, 35, 36, 37, 38, 39, 40, 41 it integrates nested sampling, Wang–Landau, Metropolis Monte Carlo, and (for suitable lattice systems) exact enumeration into a single extensible framework. This framework spans both atomistic and lattice representations with interchangeable energy calculators and move sets, enabling direct, reproducible benchmarking and comparison of interfacial phase equilibria across methods. FreeBird.jl is an active project; planned updates include superposition-enhanced nested sampling, 42 parallel tempering, 43 transition-matrix Monte Carlo, 44 geometry-aware coarse-graining (e.g., Gay–Berne 45, 46 and patchy-particle 47 models), and full GPU acceleration. Our long-term vision is to position FreeBird.jl at the core of autonomous, multi-fidelity workflows (integrating Bayesian experimental design, interface-aware machine-learning potentials, and a surface-structure database) to accelerate the discovery and design of functional interfaces.
@@ 2 Implementations
In this section, we outline the design philosophy and structure of FreeBird.jl, along with the data types it implements to enable efficient execution. We then describe how these data types support the construction of computational models for both continuous atomistic and discretized lattice systems. Finally, we summarize the configurational sampling methods currently available in FreeBird.jl, providing a brief introduction to technical details relevant to our demonstrations.
@@ 2.1.3 Research Software Development Best Practices
Documenter.jl 48 is used to automatically generate and deploy FreeBird.jl’s documentation to GitHub Pages. It collects and renders method docstrings for inclusion in the documentation, ensuring that function descriptions remain synchronized with the source code. All exported methods are required to have docstrings describing their functionality, input arguments, and return values (if any). Documentation for any function can also be accessed directly from the Julia REPL. Documenter.jl builds separate pages for stable, development, and release versions of the code, and generates a preview when a pull request is opened. The FreeBird.jl documentation also includes tutorials and examples to support new users and developers.
@@ 2.1.4 Parallelization
Julia supports parallelization via the Message Passing Interface (MPI) protocol through external packages such as MPI.jl. 49 GPU computing is also enabled by hardware-specific packages, including CUDA.jl 50, 51, 52 for NVIDIA, AMDGPU.jl 53 for AMD, oneAPI.jl 54 for Intel, and Metal.jl 55 for Apple Silicon GPUs. Additionally, vendor-neutral GPU programming models such as OpenCL.jl 56 and KernelAbstractions.jl 57 allow Julia code to run across multiple platforms from different hardware vendors.
@@ 2.2.1 Atomistic Systems with Interatomic Potentials
We use the term atomistic to describe a system composed of individual particles whose positions vary continuously in real space, confined within a simulation cell that may employ periodic boundary conditions. Configuration space is explored using single-particle displacement moves, trial translations with directions chosen uniformly at random and magnitudes drawn from a prescribed distribution. The sequence of accepted moves forms a random walk through configuration space, which is the default sampling scheme in FreeBird.jl (among others). Listing 2 illustrates how to generate a set of random atomistic configurations and store them in an AtomWalker{C} data type.
@@
As noted earlier, in AtomWalker{C}, C is a positive integer specifying the number of components (e.g., distinct chemical species or rigid subgroups) represented by the walker. In Listing 2, line 23 demonstrates how a system of five particles in a cubic box can be divided into two components, with the first component frozen. Frozen particles are excluded from subsequent single-particle displacement moves; their energy is computed once and stored, while interactions with mobile particles are still calculated on the fly, thereby reducing overall computational cost.
@@
Particles in an atomistic system interact through (typically pairwise) interatomic potentials, implemented as subtypes of AbstractPotential. For example, in Listing 3, line 2, the LJParameters, a subtype of LennardJonesParameterSets, is constructed using keyword arguments. In this example, specifying a potential cutoff at automatically shifts the potential energies so they approach 0 eV at the cutoff. This shift can be disabled by supplying the additional argument shift=false. For multi-component systems, CompositeLJParameters{C} can be used to define interactions. This data structure, a subtype of LennardJonesParameterSets, stores a matrix of LJParameters.
Lastly, Listing 3 illustrates how to construct a live set by combining a set of walkers with an LJ potential. LJAtomWalkers, a subtype of AtomWalkers (itself a subtype of AbstractLiveSet), contains AtomWalker configurations of particles interacting via LJ potentials. By default, when walkers and a potential are passed to LJAtomWalkers, their energies are automatically computed and updated. For example, the walkers in wks in line 33 of Listing 2 initially have no assigned energies; invoking LJAtomWalkers(wks, lj) in line 6 of Listing 3 evaluates those energies using the potential lj defined in line 2. This design allows any set of walkers to be paired with any potential while keeping energies current. Supplying the keyword argument assign_energy=false disables this feature, leaving energies unchanged. Many functions in FreeBird.jl follow this pattern, where expected default behavior can be overridden through additional user-provided arguments. This approach supports robust code execution and reliable user interaction while enabling developers to explore unconventional settings.
@@ 2.2.2 Lattice Systems with Discrete Hamiltonians
Lattice configurations are represented by the built-in, parameterized data type MLattice{C,G}, where C specifies the number of components (in the chemical thermodynamic sense) and G denotes the lattice geometry. MLattice{C,G} is a subtype of AbstractLattice. Listing 4 provides two examples of constructing two-dimensional lattice systems in FreeBird.jl: (1) a single-component square lattice (C=1 by default, G=SquareLattice) created using SLattice{G}, an alias for MLattice{1,G}, and (2) a two-component triangular lattice (C=2, G=TriangularLattice). In both cases, only the keyword argument components is passed to the constructor, with all other fields (lattice vectors, basis, and supercell dimensions) set to their default values but modifiable by the user. To model adsorption processes, a subset of lattice grid points can be designated as adsorption sites. Particles occupying these sites experience an on-site energy in addition to their neighboring interaction energies.
@@
The energy of a lattice can be efficiently evaluated using a discrete Hamiltonian, which specifies only the on-site energy and as many -th nearest-neighbor interactions as needed. Because a lattice is defined on a discretized grid rather than by continuous distances, a lattice gas model is considerably more efficient than an interatomic potential, where energy depends explicitly on distance. For example, Listing 5 shows the construction of a generic lattice Hamiltonian, specifying the magnitudes of the on-site, first-nearest, and second-nearest neighbor interactions, along with the units of energy. The lattice Hamiltonian is given by
where is the number of occupied sites, is the energy of an occupied site, is the nearest-neighbor interaction energy, and is the next-nearest-neighbor interaction energy. We pass these interaction values into a GenericLatticeHamiltonian{N,U}, as shown in Listing 5, where N is the number of -th nearest-neighbor interaction terms included (two in this example) and U specifies the energy units. The Hamiltonian is independent of specific lattice geometry and can therefore be applied to any lattice type. For multi-component systems, a matrix of GenericLatticeHamiltonian objects can be defined for inter- and intra-component interactions using MLatticeHamiltonian{C,N,U}, where C is the number of components.
@@ 2.3.2 Wang–Landau Sampling
The Wang–Landau (WL) algorithm is an MC method that achieves uniform sampling across a system’s energy range by iteratively refining an estimate of the density of states, . 59 This adaptive strategy enables thorough exploration of configuration space, including rugged or frustrated energy landscapes where conventional Metropolis algorithms may become trapped in metastable states. The algorithm proceeds as follows:
@@ 2.3.3 Nested Sampling
The nested sampling (NS) algorithm is an MC method in which the accessible phase space contracts stochastically, producing approximately uniform steps in , the logarithm of the remaining phase-space volume. 60, 61 This transformation converts the high-dimensional canonical partition function integral into a tractable one-dimensional sum. 62, 21, 22 The algorithm proceeds as follows:
@@ 3 Demonstrations
In this section, we demonstrate how FreeBird.jl can be used to identify the number, temperature, and nature of phase transitions across diverse systems. These span single- and multi-component compositions, spatial resolutions from continuous atomistic models to discretized lattice models, and dimensionalities in both two and three dimensions. We also benchmark three sampling methods (NS, MC, and WL) to illustrate FreeBird.jl’s versatility and to validate its predictions. The complete set of simulation parameter values is provided in the Julia code associated with each demonstration in Figures 6–9. For clarity and reproducibility, we focus on relatively simple model systems. In each case, we compare constant-volume heat capacity () curves as functions of temperature, identifying phase transitions from the peak positions, which are determined using the findmaxima() function in the Julia package Peaks.jl. For MC simulations, because runs were performed at discrete temperatures that may not coincide exactly with a transition temperature, the reported MC transition temperature corresponds to the simulated point nearest each identified phase transition. Each sampling method is executed three times to estimate the uncertainty in the resulting profiles. Although these systems contain only a small number of particles, they provide minimal yet chemically and physically meaningful representations.
@@ 3.1.1 LJ6 Cluster
We begin with the LJ6 cluster, with benchmark results available from NS. 62 We use the shifted, truncated LJ potential parameters from ref. 29 ( eV, Å, cutoff ). The system consists of six free LJ particles in a non-periodic cubic box with an edge length of 15 Å, corresponding to a number density of . We performed NS simulations with 120 walkers and a constant walker-removal rate of (Section 2.3.3), using the MCRandomWalkClone() algorithm to generate new configurations by cloning an existing walker and subsequently decorrelating it via random walks. For MC simulations, the temperature was decreased in 50 K increments from 1000 K to 50 K. For WL sampling, we employed 1000 bins spanning the energy range eV (slightly above the ground-state energy of -1.27 eV), applied an 80 % flatness criterion to the energy histogram, and used an -schedule of .
Previous studies on a lower-density LJ6 system () reported two phase transitions: a gas–solid transition at and a lower-temperature solid–solid transition at , with the latter favoring an octahedral ground state. 62 Figure 6 shows that all three methods recover these transitions, with consistent transition temperatures in units of : 0.353 (NS), 0.345 (MC), and 0.352 (WL) for the gas–solid transition; and 0.082 (NS), 0.086 (MC), and 0.101 (WL) for the solid–solid transition.
@@ 3.1.2 Binary LJ Cluster
Next, we examine the binary A1B12 cluster, for which geometry-optimized structures are available as benchmarks. 63 We reuse the shifted, truncated LJ potential from Section 3.1.1, assigning distinct well depths to represent the two components: eV for A–A interactions, eV for B–B interactions, and eV (Berthelot rule) for A–B interactions. To isolate the effect of interaction-energy asymmetry on the curve, we set the particle sizes equal ( Å). The particles are placed in a non-periodic cubic box of edge length 15.41 Å, corresponding to a number density of . Simulations were run as follows: NS simulations used 960 walkers with 400 decorrelation steps, a significant increase over LJ6 to account for the more than doubled degrees of freedom from 13 particles; MC was cooled from 400 to 100 K in 25 K increments; and WL sampled energies from 0 to -2.44 eV, slightly above the ground-state energy (-2.45 eV).
As with LJ6, the A1B12 cluster exhibits a gas–solid transition at higher temperatures and a solid–solid transition at lower temperatures. Unlike LJ6’s octahedral ground state, however, the low-temperature structure adopts an A–B icosahedral core–shell configuration. 63 Figure 7 shows that all three methods reproduce these features, with closely matching transition temperatures in units of : 0.276 (NS), 0.280 (MC), and 0.275 (WL) for the gas–solid transition, and 0.169 (NS), 0.172 (MC), and 0.167 (WL) for the solid–solid transition. Compared to LJ6, the A1B12 system shows a slightly lower gas–solid transition temperature and a markedly higher, sharper solid–solid transition.

### v3

@@ Abstract
Department of Computer Science and Engineering, McKelvey School of Engineering, Washington University in St. Louis, St. Louis, MO 63130, USA
@@ 1 Introduction
Because high-resolution in situ probes remain scarce, 10, 11 theoretical and computational approaches have become indispensable for mapping the configurational landscape of interfaces and guiding materials design. 12, 13, 14, 15 Modeling an interface under working conditions is essentially a non-equilibrium surface problem whose dynamics depend on the poorly constrained, synthesis-specific starting structure. 16, 17 To address this, researchers typically begin by establishing an equilibrium thermodynamic baseline, 18, 19 which entails a series of decisions: the level of energy evaluation (first-principles versus classical force fields) 20, 21 and the depth of configurational sampling required to obtain entropies and free energies (harmonic, quasi-harmonic, or fully anharmonic treatments; atomistic versus lattice representations). 22, 23, 24, 25
@@
Most realistic surface problems are too complex for exhaustive first-principles treatment, forcing researchers to navigate an accuracy–sampling trade-off. 18, 26 Ab initio thermodynamics, the workhorse for predicting surface phase diagrams, typically evaluates only tens to thousands of candidate structures within a practical computing budget, with selections often guided by chemical intuition. 18, 19 Structure-search algorithms such as particle-swarm, 27 Bayesian, 28, 29 and genetic 30, 31 optimizers expand the candidate pool but still assume a sharply truncated configuration space in which atoms vibrate harmonically about fixed lattice sites. 22, 24 This harmonic perspective understates the structural and compositional agility of catalytic surfaces; in many cases, comprehensive sampling with well-parameterized classical potentials can reveal low-energy motifs that sparse, high-accuracy DFT surveys may miss. 20, 21, 32, 33
@@
Although progress has been made toward systematically improvable thermodynamic sampling of surface phases, 34, 35, 32 the field still lacks a fast, reproducible benchmarking platform dedicated to interfacial structure prediction, applicable to solid–vacuum/gas systems as well as defects at solid–solid and solid–liquid boundaries. We therefore introduce FreeBird.jl, an open-source Julia package featuring a highly parallel, modular architecture that allows users to interchange system resolutions, energy models, and Monte Carlo–style algorithms without rewriting code. Distinct from existing packages, 36, 37, 38, 39, 40, 41, 42, 43, 44 it integrates nested sampling, Wang–Landau, Metropolis Monte Carlo, and (for suitable lattice systems) exact enumeration into a single extensible framework. This framework spans both atomistic and lattice representations with interchangeable energy calculators and move sets, enabling direct, reproducible benchmarking and comparison of interfacial phase equilibria across methods. FreeBird.jl is an active project; planned updates include superposition-enhanced nested sampling, 45 parallel tempering, 46 transition-matrix Monte Carlo, 47 geometry-aware coarse-graining (e.g., Gay–Berne 48, 49 and patchy-particle 50 models), and full GPU acceleration. Our long-term vision is to position FreeBird.jl at the core of autonomous, multi-fidelity workflows (integrating Bayesian experimental design, interface-aware machine-learning potentials, and a surface-structure database) to accelerate the discovery and design of functional interfaces.
@@ 2 Implementations
In this section, we outline the design philosophy and structure of FreeBird.jl, along with the data types it implements to enable efficient execution. We then describe how these data types support the construction of computational models for both continuous atomistic and discretized lattice systems. Finally, we summarize the configurational sampling methods currently available in FreeBird.jl, providing a brief introduction to technical details relevant to our demonstrations. All results reported here were obtained using FreeBird.jl release version 0.2.1.
@@ 2.1.3 Research Software Development Best Practices
Documenter.jl 51 is used to automatically generate and deploy FreeBird.jl’s documentation to GitHub Pages. It collects and renders method docstrings for inclusion in the documentation, ensuring that function descriptions remain synchronized with the source code. All exported methods are required to have docstrings describing their functionality, input arguments, and return values (if any). Documentation for any function can also be accessed directly from the Julia REPL. Documenter.jl builds separate pages for stable, development, and release versions of the code, and generates a preview when a pull request is opened. The FreeBird.jl documentation also includes tutorials and examples to support new users and developers.
@@ 2.1.4 Parallelization
Julia supports parallelization via the Message Passing Interface (MPI) protocol through external packages such as MPI.jl. 52 GPU computing is also enabled by hardware-specific packages, including CUDA.jl 53, 54, 55 for NVIDIA, AMDGPU.jl 56 for AMD, oneAPI.jl 57 for Intel, and Metal.jl 58 for Apple Silicon GPUs. Additionally, vendor-neutral GPU programming models such as OpenCL.jl 59 and KernelAbstractions.jl 60 allow Julia code to run across multiple platforms from different hardware vendors.
@@ 2.2 Models
Although the present work highlights applications to monoatomic solids, forthcoming releases will extend FreeBird.jl to molecular systems by integrating ASE calculators 61 for classical 62, 63, 64 and machine-learning 65, 66, 67 force fields and coupling with Molly.jl 68 to enable molecular dynamics–based sampling of bonds, angles, and dihedrals.
We provide code examples to illustrate the typical Julia syntax used in FreeBird.jl, but these listings are entirely optional. Skipping them will not affect the overall comprehensibility of the work.
@@ 2.2.1 Atomistic Systems with Interatomic Potentials
We use the term atomistic to describe a system composed of individual particles whose positions vary continuously in real space, confined within a simulation cell that may employ periodic boundary conditions. Configuration space is explored using single-particle displacement moves, trial translations with directions chosen uniformly at random and magnitudes drawn from a prescribed distribution. The sequence of accepted moves forms a random walk through configuration space, which is the default sampling scheme in FreeBird.jl (among others). Figure 2 illustrates how to generate a set of random atomistic configurations and store them in an AtomWalker{C} data type.
@@
As noted earlier, in AtomWalker{C}, C is a positive integer specifying the number of components (e.g., distinct chemical species or rigid subgroups) represented by the walker. In Figure 2, line 23 demonstrates how a system of five particles in a cubic box can be divided into two components, with the first component frozen. Frozen particles are excluded from subsequent single-particle displacement moves; their energy is computed once and stored, while interactions with mobile particles are still calculated on the fly, thereby reducing overall computational cost.
@@
Particles in an atomistic system interact through interatomic potentials, implemented as subtypes of AbstractPotential, currently classified as single- or multi-component and as either pairwise or many-body. For example, in Figure 3, line 2, the LJParameters (a single-component, pairwise potential) is constructed using keyword arguments. In this example, specifying a potential cutoff at automatically shifts the potential energies so they approach 0 eV at the cutoff. This shift can be disabled by supplying the additional argument shift=false. For multi-component systems, CompositeParameterSets can be used to define interactions. This data structure, a subtype of MultiComponentPotential, stores a matrix of LJParameters.
Lastly, Figure 3 illustrates how to construct a live set by combining a set of walkers with an LJ potential. LJAtomWalkers, a subtype of AtomWalkers (itself a subtype of AbstractLiveSet), contains AtomWalker configurations of particles interacting via LJ potentials. By default, when walkers and a potential are passed to LJAtomWalkers, their energies are automatically computed and updated. For example, the walkers in wks in line 33 of Figure 2 initially have no assigned energies; invoking LJAtomWalkers(wks, lj) in line 6 of Figure 3 evaluates those energies using the potential lj defined in line 2. This design allows any set of walkers to be paired with any potential while keeping energies current. Supplying the keyword argument assign_energy=false disables this feature, leaving energies unchanged. Many functions in FreeBird.jl follow this pattern, where expected default behavior can be overridden through additional user-provided arguments. This approach supports robust code execution and reliable user interaction while enabling developers to explore unconventional settings.
@@ 2.2.2 Lattice Systems with Discrete Hamiltonians
Lattice configurations are represented by the built-in, parameterized data type MLattice{C,G}, where C specifies the number of components (in the chemical thermodynamic sense) and G denotes the lattice geometry. MLattice{C,G} is a subtype of AbstractLattice. Figure 4 provides two examples of constructing two-dimensional lattice systems in FreeBird.jl: (1) a single-component square lattice (C=1 by default, G=SquareLattice) created using SLattice{G}, an alias for MLattice{1,G}, and (2) a two-component triangular lattice (C=2, G=TriangularLattice). In both cases, only the keyword argument components is passed to the constructor, with all other fields (lattice vectors, basis, and supercell dimensions) set to their default values but modifiable by the user. To model adsorption processes, a subset of lattice grid points can be designated as adsorption sites. Particles occupying these sites experience an on-site energy in addition to their neighboring interaction energies.
@@
The energy of a lattice can be efficiently evaluated using a discrete Hamiltonian, which specifies only the on-site energy and as many -th nearest-neighbor interactions as needed. Because a lattice is defined on a discretized grid rather than by continuous distances, a lattice gas model is considerably more efficient than an interatomic potential, where energy depends explicitly on distance. For example, Figure 5 shows the construction of a generic lattice Hamiltonian, specifying the magnitudes of the on-site, first-nearest, and second-nearest neighbor interactions, along with the units of energy. The lattice Hamiltonian is given by
where is the number of occupied sites, is the energy of an occupied site, is the nearest-neighbor interaction energy, and is the next-nearest-neighbor interaction energy. We pass these interaction values into a GenericLatticeHamiltonian{N,U}, as shown in Figure 5, where N is the number of -th nearest-neighbor interaction terms included (two in this example) and U specifies the energy units. The Hamiltonian is independent of specific lattice geometry and can therefore be applied to any lattice type. For multi-component systems, a matrix of GenericLatticeHamiltonian objects can be defined for inter- and intra-component interactions using MLatticeHamiltonian{C,N,U}, where C is the number of components.
@@ 2.3.2 Wang–Landau Sampling
The Wang–Landau (WL) algorithm is an MC method that achieves uniform sampling across a system’s energy range by iteratively refining an estimate of the density of states, . 70 This adaptive strategy enables thorough exploration of configuration space, including rugged or frustrated energy landscapes where conventional Metropolis algorithms may become trapped in metastable states. The algorithm proceeds as follows:
@@ 2.3.3 Nested Sampling
The nested sampling (NS) algorithm is an MC method in which the accessible phase space contracts stochastically, producing approximately uniform steps in , the logarithm of the remaining phase-space volume. 71, 72 This transformation converts the high-dimensional canonical partition function integral into a tractable one-dimensional sum. 73, 23, 24 The algorithm proceeds as follows:
@@ 3 Demonstrations
In this section, we demonstrate how FreeBird.jl can be used to identify the number, temperature, and nature of phase transitions across diverse systems. These span single- and multi-component compositions, spatial resolutions from continuous atomistic models to discretized lattice models, and dimensionalities in both two and three dimensions. We also benchmark three sampling methods (NS, MC, and WL) to illustrate FreeBird.jl’s versatility and to validate its predictions. The complete set of simulation parameter values is provided in the Julia code associated with each demonstration in Figures 6–9. For clarity and reproducibility, we focus on relatively simple model systems. In each case, we compare constant-volume heat capacity () curves as functions of temperature, identifying phase transitions from the peak positions, which are determined using the findmaxima() function in the Julia package Peaks.jl. For MC simulations, because runs were performed at discrete temperatures that may not coincide exactly with a transition temperature, the reported MC transition temperature corresponds to the simulated point nearest each identified phase transition. Each sampling method is executed three times to estimate the uncertainty in the resulting profiles. The reported values are the means of three runs, with the minimum and maximum at each temperature providing the lower and upper bounds of the uncertainty (hereafter referred to as min–max). Although these systems contain only a small number of particles, they provide minimal yet chemically and physically meaningful representations.
@@ 3.1.1 LJ6 Cluster
We begin with the LJ6 cluster, with benchmark results available from NS. 73 We use the shifted, truncated LJ potential parameters from ref. 32 ( eV, Å, cutoff ). The system consists of six free LJ particles in a non-periodic cubic box with an edge length of 15 Å, corresponding to a number density of . We performed NS simulations with 120 walkers and a constant walker-removal rate of (Section 2.3.3), using the MCRandomWalkClone() algorithm to generate new configurations by cloning an existing walker and subsequently decorrelating it via random walks. For MC simulations, the temperature was decreased in 50 K increments from 1000 K to 50 K. For WL sampling, we employed 1000 bins spanning the energy range eV (slightly above the ground-state energy of -1.27 eV), applied an 80 % flatness criterion to the energy histogram, and used an -schedule of .
Previous studies on a lower-density LJ6 system () reported two phase transitions: a gas–solid transition at and a lower-temperature solid–solid transition at , with the latter favoring an octahedral ground state. 73 Figure 6 shows that all three methods recover these transitions, with consistent transition temperatures in units of : 0.353 (NS), 0.345 (MC), and 0.352 (WL) for the gas–solid transition; and 0.082 (NS), 0.086 (MC), and 0.101 (WL) for the solid–solid transition. This low-temperature transition has also been described in previous studies of small Lennard-Jones clusters as a melting from a solid to a liquid-like state, with reported transition temperatures in good agreement. 73, 74, 75
@@ 3.1.2 Binary LJ Cluster
Next, we examine the binary A1B12 cluster, for which geometry-optimized structures are available as benchmarks. 76 We reuse the shifted, truncated LJ potential from Section 3.1.1, assigning distinct well depths to represent the two components: eV for A–A interactions, eV for B–B interactions, and eV (Berthelot rule) for A–B interactions. To isolate the effect of interaction-energy asymmetry on the curve, we set the particle sizes equal ( Å). The particles are placed in a non-periodic cubic box of edge length 15.41 Å, corresponding to a number density of . Simulations were run as follows: NS simulations used 960 walkers with 400 decorrelation steps, a significant increase over LJ6 to account for the more than doubled degrees of freedom from 13 particles; MC was cooled from 400 to 100 K in 25 K increments; and WL sampled energies from 0 to -2.44 eV, slightly above the ground-state energy (-2.45 eV).
As with LJ6, the A1B12 cluster exhibits a gas–solid transition at higher temperatures and a solid–solid transition at lower temperatures. Unlike LJ6’s octahedral ground state, however, the low-temperature structure adopts an A–B icosahedral core–shell configuration. 76 Figure 7 shows that all three methods reproduce these features, with closely matching transition temperatures in units of : 0.276 (NS), 0.280 (MC), and 0.275 (WL) for the gas–solid transition, and 0.169 (NS), 0.172 (MC), and 0.167 (WL) for the solid–solid transition. Compared to LJ6, the A1B12 system shows a slightly lower gas–solid transition temperature and a markedly higher, sharper solid–solid transition. The ground-state structure corresponds to the minimum-energy configuration in the binary Lennard-Jones database, 76 and the double-peaked curve and associated phase changes are consistent with descriptions from early work using argon- and neon-like Lennard-Jones parameters.77